Question3 A parallel plate capacitor is formed from two circular plates, each of radius 4.60 cm, separated by 2.30 mm. The plates are charged to 11.6 nC. A proton is shot from the negative plate to the positive. What launch speed must the proton have to just barely reach the positive plate? 2.08x105 m/s 4.34x106 m/s 8.69x106 m/s 2.95x105 m/s
Answer #1

Electric field, E = Q/ Ae0

E = (11.6 x 10^-9)/(pi x 0.046^2 x 8.85 x 10^-12)

E = 1.97 x 10^5 N/C

F = ma = qE

Acceleration, a = (1.6 x 10^-19 x 1.97 x 10^5)/(1.67 x 10^-27) = 1.89 x 10^13 m/s^2

V^2 = u^2 + 2as

V^2 = 0 + 2 x 1.89 x 10^13 x 0.0023

V = 2.95 x 10^5 m/s

Comment in case any doubt please rate my answer ..